OpenStudy (anonymous):

What value of x makes the equation true? 7 over 3= x over 6

OpenStudy (zzr0ck3r):

\[\frac{7}{3}= \frac{x}{6}?\]

OpenStudy (anonymous):

yes^

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (bahrom7893):

multiply both sides by 6. What do you get?

OpenStudy (anonymous):

7 and 3?

OpenStudy (bahrom7893):

no everything: \[6*\frac{7}{3}=\frac{x}{6}*6\]

OpenStudy (anonymous):

oh ok thanks

OpenStudy (anonymous):

?-? but whats x?

OpenStudy (bahrom7893):

well what do you get after you simplify? \[2*7=x\]\[x=14\]

OpenStudy (anonymous):

14!? :)

OpenStudy (bahrom7893):

yea

OpenStudy (anonymous):

thanks that's on of these answers :D